# OpenLKA: An Open Dataset of Lane Keeping Assist from Recent Car Models under Real-world Driving Conditions

### v1

@@ V Dataset Applications
    LKA benchmarking and real-world evaluation: This dataset contains real-world data of LKA systems of different vehicle makes in different scenarios, which we can use to quantify the level of LKA in different car makes. In addition, we can also use it to quantify the most challenging scenarios of LKA and analyze the causes, thus providing guidance for LKA development and traffic safety prevention.
@@
    Enhanced LKA Design Through multimodal Data and Foundation Models: The OpenLKA dataset provides a comprehensive foundation for designing more robust LKA systems by offering diverse variables related to lane line detection and vehicle control dynamics. We also have end-to-end model outputs from OpenPilot, serving as a valuable baseline for comparative analysis. To demonstrate the dataset’s utility for advanced driving automation research, we conducted experiments with multimodal language models as shown in Fig. 12. We sampled representative driving scenarios from OpenLKA, preparing synchronized video frames and CAN bus data as inputs, with human drivers’ subsequent steering angles as labels. Through Supervised Fine-Tuning (SFT) using ChatGPT-4o’s API on these multimodal inputs, we observed significant improvements in steering prediction accuracy. These results highlight the considerable potential of leveraging multimodal foundation models for automated driving decision-making based on OpenLKA.
@@
    Infrastructure Readiness Assessment and Roadway Safety Planning: OpenLKA enables transportation infrastructure stakeholders to evaluate how current road conditions interact with LKA system performance. By identifying segments where LKA systems frequently disengage or exhibit poor lane-keeping—especially under conditions such as degraded lane markings, sharp curves, or inconsistent geometry—stakeholders can assess the roadway’s readiness for assisted driving technologies. The dataset facilitates targeted interventions, such as pavement re-striping, lane marking restoration, or geometric redesign, to reduce LKA failure risk and improve overall lane-keeping safety at the infrastructure level.
@@
    Analysis of LKA Failure to Driver Takeover Dynamics: When LKA systems fail or exhibit significant lane deviation, drivers often have little to no time to react, posing safety risks. The OpenLKA dataset provides a comprehensive record of the entire process from LKA failure or large lane deviation to driver takeover, capturing multimodal signals such as CAN data (e.g., steering angle), video footage, and VLM-annotated environmental contexts (e.g., faded markings, sharp curves). This enables researchers to analyze LKA failure modes, quantify the time window from failure to takeover, and model driver reaction patterns under diverse conditions [28]. Such insights can inform the development of proactive warning systems that predict LKA instability and alert drivers in advance, enhancing overall safety.
Figure 12: Steering angle prediction using GPT-4o: The original model predicts 6.85°, while fine-tuning with OpenLKA data yields 7.52°, matching the human driver’s actual action. (7.52°).
@@ VI. Conclusion
The OpenLKA dataset pioneers a uniquely comprehensive, multimodal, and real-world foundation for advancing LKA system research. Its primary strength lies in its integration of synchronized CAN bus, high-definition dashcam videos, detailed Openpilot outputs for precise lane-keeping metrics, and extensive contextual annotations via VLM. Furthermore, OpenLKA uniquely leverages continuous data contributions from a large, global driving community, ensuring ongoing dataset expansion and diversity.
However, it lacks data on trailing vehicles, which is critical for comprehensive traffic studies. Future work will integrate radar and LiDAR to enhance data collection, and leverage the dataset to evaluate LKA performance in real-world scenarios, propose improvements for in-vehicle systems and roadside infrastructure, and develop human-centered warning systems for smoother ADAS takeovers.

### v2

@@ V Discussion: Dataset Applications
    LKA benchmarking and real-world evaluation:
    This real‑world dataset spans multiple vehicle brands and operating conditions, allowing us to (i) benchmark Lane‑Keeping Assist (LKA) performance by make and (ii) isolate the scenarios most prone to failure. Figure.11 ranks the leading single factors—such as faded lane markings, poor pavement‑lane contrast, and sharp curvature—and their high‑impact combinations, revealing that visibility‑related degradations account for more than 60% of recorded disengagements. These findings offer concrete guidance for both OEM algorithm refinement and infrastructure interventions, which we will elaborate in forthcoming work.
@@
    Multimodal Foundation Models & VLA: Leveraging frame‑aligned video, CAN dynamics, and OP trajectories, OpenLKA is a turnkey test bed for Vision–Language–Action (VLA) research[29]. Our preliminary experiments with multimodal foundation models (Fig. 12) show promising gains in steering prediction, and we will continue exploring this avenue for next‑generation LKA and VLA policy design.
@@
    Infrastructure Readiness Assessment and Roadway Safety Planning: OpenLKA lets infrastructure stakeholders diagnose where and why LKA systems falter. Our current work already quantifies how roadway factors—e.g., the relationship between lane deviation and curvature (in Fig. 13) shape LKA performance, and fuller results will appear in a follow‑up paper. Building on these analyses, practitioners can pinpoint segments with frequent disengagements, then direct counter‑measures such as pavement re‑striping, lane‑marking renewal, or geometric redesign to boost assisted‑driving safety.
@@
    Analysis of LKA Failure to Driver Takeover Dynamics: When LKA systems fail or exhibit significant lane deviation, drivers often have little to no time to react, posing safety risks. OpenLKA provides a comprehensive record of the entire process from LKA failure or large lane deviation to driver takeover, capturing multimodal signals. This enables researchers to analyze LKA failure modes, quantify the time window from failure to takeover, and model driver reaction patterns under diverse conditions [30]. Such insights can inform the development of proactive warning systems that predict LKA instability and alert drivers in advance, enhancing human-machine interaction safety.
Figure 11: Dominant causes of LKA failure extracted from OpenLKA: (a) share of single‑factor failures; (b) top multi‑factor combinations and their prevalence.
Figure 12: Steering angle prediction using GPT-4o: The original model predicts 6.85°, while fine-tuning with OpenLKA data yields 7.52°, matching the human driver’s actual action.
Figure 13: Quantitative analysis of LKA’s lane deviation and road curvature; the lane deviation increases with increasing road curvature, revealing an approximately linear relationship.
@@ VI. Conclusion
The OpenLKA dataset pioneers a uniquely comprehensive, multimodal, and real-world foundation for advancing LKA system research. Its primary strength lies in its ground truth of decoded CAN signals that reveal real-world LKA performance, supplemented by synchronized dashcam videos, and extensive contextual annotations via VLM. Furthermore, OpenLKA uniquely leverages continuous data contributions from a large, global driving community, ensuring ongoing dataset expansion and diversity.
The dataset still lacks heavy-duty vehicles, whose steering challenges and lane departure consequences can be even more critical than passenger cars. Future work on this dataset will focus on extension to commercial trucks.